OpenStudy (anonymous):

(08.01, 08.02, 08.03, 08.05, 08.06) Part A: Factor 3x2y2 - xy2 - 2y2. Show your work. (3 points) Part B: Factor x2 + 6x + 9. Show your work. (3 points) Part C: Factor x2 - 25. Show your work. (2 points) Part D: Do the polynomials in parts A, B, and C have a common factor? Explain your answer. (2 points)

OpenStudy (anonymous):

\[3x^2y^2-xy^2-2y^2=y^2(3x^2-x-2)=y^2(x-1)(3x+2)\]

OpenStudy (anonymous):

\[x^2+6x+9=(x+3)^2\]

OpenStudy (anonymous):

\[x^2-25=(x-5)(x+5)\]

OpenStudy (anonymous):

they do not have a common factor because we can see from the answers, there are no two factors are alike.